OpenStudy (anonymous):

Somemone Please HELP AND DO THIS GIVING FAN AND MEDAL @ganeshie8 @Compassionate @bohotness @PurplePanda312

OpenStudy (anonymous):

Create the price function for Stock D. It should have the second lowest starting price and the second highest rate of growth. Using complete sentences, justify that your function meets these requirements. Graph the price function for Stock D. Explain the key features of the graph with complete sentences. A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x). Gordon wants to fully understand what kind of changes could affect the money he is investing. Using complete sentences, explain the effect of the following transformations. Graph them and f(x) on the same coordinate plane. f(x + 2) f(–x) f(x) + 3

OpenStudy (anonymous):

@AnswerMyQuestions @pooja195 @Nurali

pooja195 (pooja195):

@iGreen.

OpenStudy (anonymous):

@iGreen.

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

Your printscreen doesn't show everything..it's too much to the left..

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

It only shows Stocks A, B, and C on the table..D is not shown.

OpenStudy (anonymous):

no actually i am just showing you the bar the questions are already here

OpenStudy (anonymous):

Yes, but I need the entire table..

OpenStudy (anonymous):

thats all of it @iGreen.

OpenStudy (anonymous):

Can you scroll to the right?

OpenStudy (anonymous):

Oh, okay

OpenStudy (anonymous):

i promise

OpenStudy (anonymous):

Okay, nvm, I thought it would be on the table but it's not supposed to..my bad.

OpenStudy (anonymous):

thats ok

OpenStudy (anonymous):

i gtg srry i have to go to school (regular school not virtual)sorry if u can con.please do

OpenStudy (anonymous):

Okay, so we can write equations for Stocks B and C. For C, it says the starting price is 30, and it increases by 4%. So we write: \(f(x) = 30(1.04)^x\) For B, it is actually decreasing by 2%, and the starting cost is 22: \(f(x) = 22(0.98)^x\) So we have: Stock A: \(f(x) = 25(1.08)^x\) Starting price is $25, and it increases by 8%. Stock B: \(f(x) = 22(0.98)^x\) Starting price is $22, and it decreases by 2%. Stock C: \(f(x) = 30(1.04)^x\) Starting price is $30, and it increases by 4%. Stock D: It needs to have the 2nd lowest starting price and highest rate of growth. Let's say the starting price is 24, and the rate of growth is 10%. So we write: \(f(x) = 24(1.10)^x\)

OpenStudy (anonymous):

Okay, now we have to graph Stock D. Here's the graph (in the attachment)

OpenStudy (anonymous):

Key features: The y-intercept is at 24, and it increases.

OpenStudy (anonymous):

@heymydearfriend Are you there?

pooja195 (pooja195):

@heymydearfriend please do not simply look for answers igreen is helping you so you need to participate also hes done a spectacular job explaining things reward woth a medal by pressing best respone thank you! :)

OpenStudy (anonymous):

He/She is gone

OpenStudy (anonymous):

srry @igreen i had to go to school

OpenStudy (anonymous):

may you help me again this time i am here

OpenStudy (anonymous):

Oh, okay

OpenStudy (anonymous):

what were saying in the beginning

OpenStudy (anonymous):

"A recent drop in sales has affected Stock D with the function g(x) = –6. Explain to Gordon how Stock D’s new price function, f(x) + g(x), will be created. Graph f(x) + g(x)." Okay, so now we have to add the two functions together. \(f(x) = 24(1.10)^x + -6\) Or \(f(x) = 24(1.10)^x - 6\) That's our new function.

OpenStudy (anonymous):

Just look up and read what I said..right now I'll answer the rest of our questions.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@igreen may you help me w, something else before you help me w. this i need to finish to first

OpenStudy (anonymous):

@iGreen.